### v1

@@ Abstract
With the rapid development of machine learning, numerical discretization methods based on deep neural networks have been widely used in many fields, especially in solving high-dimensional problems where traditional methods face bottlenecks. However, for low-dimensional problems, existing machine learning methods are not as accurate and efficient as traditional methods. In this paper, we propose a subspace method based on neural networks for eigenvalue problems with high accuracy and low cost. Our basic idea is to use neural network based basis functions to span a subspace, then calculate the parameters of the neural network based basis functions through appropriate training, and finally calculate the Galerkin projection of the eigenvalue problem onto the subspace and obtain an approximate solution. In addition, we reduce the dimension of the subspace by applying some dimensionality reduction technique, which can improve the accuracy of the approximate solution in further. Numerical experiments show that we can obtain approximate eigenvalues with accuracy of $10^{-11}$ but with less then 400 epochs, which is significantly superior to other existing neural network based methods.
@@
subspace, basis function, eigenvalue problem, multi-eigenpairs, neural networks, high-accuracy
@@ 1 Introduction
With the rapid development of machine learning, many numerical discretization methods for partial differential equations (PDEs) based on deep neural networks have received increasing attention, such as the physical-information neural networks (PINN) [35], the deep Galerkin method (DGM) [39], the deep Ritz method (DRM) [50], the deep Nitsche method (DNM) [32], and the weak adversarial networks (WAN) [52]. By these methods, the solution of the partial differential equation is represented by a neural network directly and the optimal parameters of the neural network is obtained by minimizing a loss function. The main difference between these methods lies in the choice of the loss function. For instance, PINN and DGM use the $L^{2}$ norm of the residual in the strong form as the loss function, while DRM and WAN use the energy functional and the operator norm in the weak form as the loss function, respectively. In contrast, DNM uses a loss function based on Nitsche’s energy formulation to handle essential boundary conditions.
@@
Due to the powerful approximation capabilities of neural networks, the aforementioned methods are not only used to deal with source problems but also to handle eigenvalue problems. For instance, in [22], a wave function ansatz based on deep neural networks is introduced to solve the many-electron Schrödinger equation. In [23, 24], various wave function ansatzes in the form of a Slater determinant are also introduced for the Schrödinger equation. In [28], a neural network based multistate solver is proposed for finding the first $k$ eigenpairs of the Schrödinger equation simultaneously. In [21], a diffusion Monte Carlo-like method is proposed for handling high-dimensional eigenvalue problems. In [27], a semigroup method is provided for solving high-dimensional elliptic partial differential equations using neural networks and is extended to a primal-dual algorithm for solving eigenvalue problems. In [29], a type of tensor neural network based on the DECOMP/PARAFAC (CP) tensor decomposition is proposed, along with an efficient numerical integration scheme for tensor-type functions. Theoretical analyses and applications of tensor neural networks are provided in [44] and [43, 45], respectively. For example, tensor neural networks have been used to solve the the first eigenvalue and the corresponding eigenfunction of Schrödinger equation in [43] and the first $k$ eigenpairs of high-dimensional eigenvalue problems in [45]. Additionally, in [51], a machine learning method is proposed for computing eigenvalues and eigenfunctions of the Schrödinger equation, which employs trial functions satisfying the homogeneous boundary conditions to avoid the penalty of the boundary, and an explicit convergence rate of the generalization error also is provided in the article.
@@
While deep neural network based methods have made strides in solving high-dimensional problems where traditional methods fail, they face challenges in solving low-dimensional problems due to the lack of algorithms to train neural networks efficiently. As a result, they lag behind traditional methods in accuracy and efficiency, which prevent most deep neural network based methods from being not only usable but also efficient.
@@
In recent years, some emerging methods based on shallow neural networks have received increasing attention, such as the extreme learning machine (ELM) [25] and the random feature method (RFM) [10, 13]. In these methods, all parameters of the neural network except for the output layer are randomly generated and fixed, which simplifies the optimization problem into a linear system. By using the least-squares method to solve the linear system, the expensive training process can be carried out at a low cost with even higher accuracy. These methods and their variants [17, 38, 40] have been widely used for solving partial differential equations [10, 12, 30, 46]. However, as so far, there is little application of these methods for solving eigenvalue problems.
@@
From another perspective, the methods above fundamentally offer a method to construct basis functions using neural networks. However, as the basis functions constructed by neural networks are sensitive to the choice of hyperparameters, improper choices can significantly impact the accuracy of the approximate solution. Randomly generating these hyperparameters is proved to be inefficient for some problems, which means that the search for more reasonable ways to select and update hyperparameters is a worthwhile pursuit.
@@
To address the impact of hyperparameter selection on solutions, a modified batch intrinsic plasticity method for pre-training the randomly generated hyperparameters in ELM is proposed in [18]. A differential evolution algorithm is employed to compute the optimal or near-optimal parameter $R_{m}$, where the parameters of the neural network are set to random values generated in interval $[-R_{m},R_{m}]$ in [19]. Additionally, a better way to randomly select hyperparameters based on some geometric explanations is provided in [53]. There are also some methods that couple the updating of hyperparameters with the solution of the linear system. An alternative iteration between gradient descent to update the hyperparameters and least squares to solve the linear system is adopted in [2, 14].
@@
Recently, a subspace method based on neural networks (SNN) for partial differential equations in strong form has been proposed in [49], which provides a deterministic way to obtain the appropriate hyperparameters. The main idea of the method is to use neural network based functions as basis functions to span a subspace, then obtain the parameters of the neural network based functions by minimizing a loss function and subsequently find the approximate solution in the subspace. Unlike ELM-like methods and random feature methods, parameters of the neural network in SNN are not randomly generated; instead, they are obtained by minimizing some loss function. In comparison to classical deep neural network based discretization methods, the training process in the SNN focuses on finding better basis functions, which requires relatively lower accuracy in the training process and often needs significantly fewer training epochs. Consequently, SNN can achieve high accuracy with low cost in training for solving partial differential equations. Furthermore, a weak form version of SNN has also been proposed in [33].
@@
From another perspective, the training process of SNN essentially offers a deterministic approach to generate a basis based on neural networks. The solving process of SNN, which calculate the Galerkin projection of the original problem onto the subspace spanned by the basis functions obtained from the training process, can be viewed as a special case of finite dimensional approximation commonly employed in traditional methods. The difference lies in the selection of basis functions: traditional methods like the finite element method employs piecewise polynomials as basis functions, spectral method employs orthogonal polynomials as basis functions, whereas SNN employs appropriately trained neural network based functions as basis functions.
@@
In this paper, motivated by the work in [49], we first propose a subspace method based on neural networks for eigenvalue problems and then an improved version. In the improved version, we add a dimensionality reduction step after obtaining the subspace spanned by the neural network based basis functions. Our method mainly consist of four steps. First, we specify neural networks architecture and appropriately modify basis functions as needed for the specific problem. Next, we obtain the parameters of the basis functions by minimizing some loss function. Then, we reduce the dimension of the subspace spanned by the basis functions by some dimensionality reduction technique. Finally, we calculate the Galerkin projection of the original eigenvalue problem onto the reduced subspace and obtain an approximate solution.
@@
We apply our method to solution of some typical eigenvalue problems. In those examples, our algorithm can achieve very high accuracy with few training epochs, which demonstrates the advantages of our new algorithm in both computational cost and accuracy.
@@
The subsequent parts of our paper are organized as follows. In Section 2, we provide some preliminary knowledge, including some notation and classical results for a linear elliptic eigenvalue problem, and an briefly introduction of the subspace method based on neural networks for some source problems. In Sections 3, we introduce the general framework of the subspace method based on the neural networks for eigenvalue problems and the details for each step. In Section 4, we introduce some dimensionality reduction technique to reduce the dimension of the subspace spanned by the neural network based functions, and obtain the improved version of SNN for eigenvalue problems. In Section 5, we provide some numerical results to show the efficiency of our new method. Finally, we conclude with a brief conclusion.
@@ 2 Preliminaries
Before introducing our method, we first give some background knowledge, including some basic results of a elliptic eigenvalue problem and some briefly introduction to the subspace method based on neural networks for partial differential equations proposed in [49].
@@ 2.1 A linear elliptic eigenvalue problem and its finite dimensional discretization
[…]
@@
[…]
@@
[…]
@@
Suppose $V_{h}=\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$, where $\varphi_{1},\cdots,\varphi_{M}$ are basis functions of the finite dimensional subspace $V_{h}$. Substituting $u_{h}=\sum_{j=1}^{M}w^{(j)}\varphi_{j}\ \text{and choosing}\ v_{h}\ \text{to be}\ \varphi_{i}$ into (2), we then have:
@@
Denote $A$ the stiffness matrix and $B$ the mass matrix, that is, $A=(a(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M},B=(b(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M}$. The eigenvalue problem (1) can be formulated as the following matrix eigenvalue problem:
$Aw=\lambda Bw,$ (3)
where $w=(w^{(1)},\cdots,w^{(M)})^{T}\in\mathbb{R}^{M}.$
By solving the matrix eigenvalue problem (3), an approximate solution to the eigenvalue problem (1) can be obtained.
@@ 2.2 Subspace method based on neural networks
The neural networks architecture used in [49] is comprised of four parts: an input layer, several hidden layers, a subspace layer and an output layer. For simplicity, we take the case of single output as an example to show the mathematical expression of the neural network. The expression of multiple outputs can be obtained similarly.
Denote $L$ the number of hidden layers, $n_{0}=d$ the number of neurons in the input layer, $n_{1},\cdots,n_{L}$ the number of neurons in each hidden layer, $n_{L+1}=M$ the number of neurons in the subspace layer, and $n_{L+2}=1$ the number of neurons in the output layer. The neural network used in [49] can be expressed as follows:
@@
Here, $W_{l}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $\bm{b}_{l}\in\mathbb{R}^{n_{l}}$ are the weights and biases of the corresponding layer, respectively, $\bm{x}\in\mathbb{R}^{d}$ is the input, $\varphi=(\varphi_{1},\cdots,\varphi_{M})^{T}$ and $w=(w^{(1)},\cdots,w^{(M)})^{T}$, $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is an activate function which can be extended to $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ as follows:
$\sigma(\bm{x})=\begin{bmatrix}\sigma(x_{1})\\ \vdots\\ \sigma(x_{n})\end{bmatrix}\ \text{for any}\ \bm{x}=(x_{1},\cdots,x_{n})^{T}\ \in\mathbb{R}^{n}.$
Denote $\theta=\{W_{1},\cdots,W_{L+1},\bm{b}_{1},\cdots,\bm{b}_{L+1}\}\in\Theta:=\mathbb{R}^{n_{1}\times n_{0}}\times\cdots\times\mathbb{R}^{n_{L+1}\times n_{L}}\times\mathbb{R}^{n_{1}}\times\cdots\times\mathbb{R}^{n_{L+1}}$ the parameters in the neural network, then $u(\bm{x};\theta,w)$ is the output of $\bm{x}$ with respect to parameters $\theta$ and $w$.
Next, we will review the general framework of the subspace method based on neural networks for the partial differential equations proposed in [49].
Firstly, the neural networks architecture introduced above is adopted, and $\theta$ and $w$ are initialized. Next, fix the coefficients $w$, and update the subspace basis functions, i.e., update parameters $\theta$ by minimizing a loss function, aiming to obtain basis functions that are better than the initial ones through training. After training, calculate the Galerkin projection of the original problem onto the subspace spanned by trained basis functions $\varphi_{1},\cdots,\varphi_{M}$ and obtain an approximate solution to the partial differential equation. The above process is summarized as a general framework for solving the partial differential equations in [49]. We also list it here for the completeness, see Algorithm (1) for the details.
Algorithm 1 General framework of SNN for PDEs
  1. Initialize the neural networks architecture (4), generate randomly $\theta$, and give $w$.
  2. Update parameters $\theta$ by minimizing a loss function, and obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$.
  3. Calculate the Galerkin projection of the PDE onto the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$ and obtain an approximate solution.

### v2

@@ Abstract
In this paper, we propose a subspace method based on neural networks for eigenvalue problems with high accuracy and low cost. We first construct a neural network-based orthogonal basis by some deep learning method and dimensionality reduction technique, and then calculate the Galerkin projection of the eigenvalue problem onto the subspace spanned by the orthogonal basis and obtain an approximate solution. Numerical experiments show that we can obtain approximate eigenvalues and eigenfunctions with very high accuracy but low cost.
@@
eigenvalue problems, subspace, method neural networks orthogonal basis, Galerkin projection
@@ 1 Introduction
With the rapid development of machine learning, many numerical discretization methods for partial differential equations (PDEs) based on deep neural networks have received increasing attention, such as the physical-information neural networks (PINN) [40], the deep Galerkin method (DGM) [47], the deep Ritz method (DRM) [58], and the weak adversarial networks (WAN) [60]. By these methods, the solution of a partial differential equation is represented by neural networks directly and the optimal parameters of neural networks are obtained by minimizing some loss function. The main difference between these methods lies in the choice of the loss function. For instance, PINN and DGM use the $L^{2}$ norm of the residual in the strong form as the loss function, while DRM and WAN use the energy functional and the operator norm in the weak form as the loss function, respectively.
@@
Due to the powerful approximation capabilities of neural networks [4, 18, 25, 26, 30, 37, 38, 44, 45, 46], the aforementioned methods are used to deal with many different kinds of partial differential equations, including not only some source problems but also some eigenvalue problems. For instance, in [24], a wave function ansatz based on deep neural networks is introduced to solve the many-electron Schrödinger equation. In [Hermann_Schätzle_Noé_2020, Hermann_Spencer_Choo_Mezzacapo_Foulkes_Pfau_Carleo_Noé_2023], various wave function ansatzes in the form of a Slater determinant are also introduced for the Schrödinger equation. In [32], a neural network-based multistate solver is proposed for finding the first $k$ eigenpairs of the Schrödinger equation simultaneously. In [23], a diffusion Monte Carlo-like method is proposed for handling high-dimensional eigenvalue problems. In [31], a semigroup method is provided for solving high-dimensional elliptic partial differential equations using neural networks and is extended to a primal-dual algorithm for solving eigenvalue problems. In [33], a type of tensor neural network-based on the DECOMP/PARAFAC (CP) tensor decomposition is proposed, along with an efficient numerical integration scheme for tensor-type functions. Theoretical analyses and applications of tensor neural networks are provided in [52] and [51, 53], respectively. For example, tensor neural networks have been used to solve the smallest eigenvalue and the corresponding eigenfunction of Schrödinger equation in [51] and the first $k$ eigenpairs of high-dimensional eigenvalue problems in [53]. Additionally, in [59], a machine learning method is proposed for computing eigenvalues and eigenfunctions of the Schrödinger equation, and an explicit convergence rate of the generalization error is also provided in the article.
@@
While deep neural network-based methods have made strides in solving high dimensional problems where classical methods fail, they face challenges in solving low-dimensional problems due to the lack of algorithms to train neural networks efficiently. As a result, they lag behind classical methods in accuracy and efficiency, which prevent most deep neural network-based methods from being not only usable but also efficient.
@@
In recent years, some emerging methods based on shallow neural networks have received increasing attention, such as the extreme learning machine (ELM) [27] and the random feature method (RFM) [11, 14]. In these methods, all parameters of neural networks except for the output layer are randomly generated and fixed, which simplifies the optimization problem into a linear system. By using the least-squares method to solve the linear system, the expensive training process can be carried out at a low cost with even higher accuracy. These methods and their variants [20, 43, 49] have been widely used for solving partial differential equations [11, 13, 34, 54]. However, so far, there is little application of these methods for solving eigenvalue problems.
@@
From another perspective, the methods above fundamentally offer a method to construct basis functions using neural networks. However, as the basis functions constructed by neural networks are sensitive to the choice of hyperparameters, improper choices can significantly impact the accuracy of the approximate solution. Randomly generating these hyperparameters may be inefficient for some problems, which means that the search for more reasonable ways to select and update hyperparameters is a worthwhile pursuit.
@@
To address the impact of hyperparameter selection on solutions, a modified batch intrinsic plasticity method for pre-training the randomly generated hyperparameters in ELM is proposed in [19]. A differential evolution algorithm is employed to compute the optimal or near-optimal parameter $R_{m}$, where the parameters of the neural networks are initialized with random values generated in the interval $R_{m},R_{m}]$ in [21]. Additionally, a better way to randomly select hyperparameters based on some geometric explanations is provided in [61].
@@
There are also some deterministic algorithms to calculate hyperparameters. An alternating iteration process is introduced in [2, 15], which combines gradient descent for updating the hyperparameters and least squares for solving the linear system. A subspace method based on neural networks (SNN) for partial differential equations has been proposed in [57]. The main idea of this method is to train neural network-based functions by minimizing a loss function, and then use least squares to find the approximation of the original problem in the subspace spanned by those trained functions, with complete decoupling between training the subspace and solving the linear system. Unlike ELM-like methods and random feature methods, the parameters of the neural network-based functions in SNN are not randomly generated; instead, they are obtained by minimizing some loss function. In comparison to classical deep neural network-based discretization methods, the training process in the SNN focuses on finding better basis functions, which requires relatively lower accuracy in the training process and often needs significantly fewer training epochs. Compared with methods updating hyperparameters by alternating iteration, the updating of hyperparameters and solution of the linear system in SNN is totally decoupled, which greatly reduces the computational cost. Consequently, SNN can achieve high accuracy with low cost in training for solving partial differential equations. Furthermore, a weak form version of SNN has also been proposed in [36].
@@
In another view, the training process of SNN essentially offers a deterministic approach to update the parameters of the neural network-based functions. The solving process of SNN, which finds an approximate solution in the subspace spanned by the trained functions, can be viewed as a kind of finite dimensional approximation commonly employed in classical methods. The difference lies in the selection of basis functions: classical methods like the finite element method employ piecewise polynomials as basis functions, spectral method employs orthogonal polynomials as basis functions, whereas SNN employs trained neural network-based functions as basis functions.
@@
In this paper, motivated by the idea of using deep learning method to train the basis functions introduced in [57], we propose a subspace method based on neural networks for eigenvalue problems. Besides using the deep learning method to train some neural network-based functions, we further introduce some dimensionality reduction technique to the trained functions and obtain some neural network-based orthogonal basis. Our method mainly consists of two steps. First, we construct a neural network-based orthogonal basis by some deep learning method and dimensionality reduction technique. Then, we use some classical method, e.g., the Galerkin method, to discretize the eigenvalue problem in the subspace spanned by the orthogonal basis, and then obtain the approximate solution.
@@
We apply our method to the solution of some typical eigenvalue problems. The numerical results show that our algorithm can achieve very high accuracy with few training epochs, which demonstrates the advantages of our new algorithm in both computational cost and accuracy.
@@
The subsequent parts of our paper are organized as follows. In Section 2, we provide some preliminary knowledge, including some notation and classical results for a linear elliptic eigenvalue problem, and a brief introduction of the subspace method based on neural networks for some source problems. In Sections 3, we introduce the general framework of our subspace method based on the neural networks for eigenvalue problems and the details for each step,especially on how to construct neural network-based orthogonal basis. In Section 4, we provide some numerical results to show the efficiency of our new method. Finally, we conclude with a brief conclusion.
@@ 2 Preliminaries
Before introducing our method, we first give some preliminary knowledge, including some basic results of a elliptic eigenvalue problem and a brief introduction to the subspace method based on neural networks for partial differential equations proposed in [57].
@@ 2.1 A linear elliptic eigenvalue problem and its finite dimensional discretization
Let ${\rm\Omega}\subset\mathbb{R}^{d}(d\geq 1)$ be a polygonal domain. We shall use the standard notation for Sobolev spaces $W^{s,p}(\rm\Omega)$ and their associated norms (see, e.g. [1]). We denote $H_{0}^{1}({\rm\Omega})=\{v\in H^{1}({\rm\Omega}):v|_{\partial{\rm\Omega}}=0\}$, $(\cdot,\cdot)$ be the standard $L^{2}$ inner product and $\langle\cdot,\cdot\rangle$ be the standard Euclidean inner product. Additionally, in this paper, we will use the letter $C$ to denote a generic positive constant which may stand for different values at its different occurrences. For convenience, following [56], the symbols $\lesssim$ are used in this paper, where $x\lesssim y$ means that $x\leq Cy$ for some constant $C$ that is independent of the spatial discretization.
@@
Here, $(V,\|\cdot\|)$ is a Hilbert space with inner product ($\cdot$$\cdot$) and norm $\|\cdot\|$. The bilinear forms $a(\cdot,\cdot)$ and $b(\cdot,\cdot)$ on $V\times V$ are symmetric and satisfy
@@
Suppose the norm $\|\cdot\|_{b}$ is compact with respect to $\|\cdot\|$ ([7]), i.e., from any sequence which is bounded in $\|\cdot\|$, one can extract a subsequence which is Cauchy in $\|\cdot\|_{b}$.
@@
Suppose $V_{h}=\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$, where $\{\varphi_{i}\}_{i=1}^{M}$ is a basis of the finite dimensional subspace $V_{h}$. Substituting $u_{h}=\sum_{j=1}^{M}w^{(j)}\varphi_{j}\ \text{and choosing}\ v_{h}\ \text{to be}\ \varphi_{i}$ into (2), we then have:
@@
Denote $A$ the stiffness matrix and $B$ the mass matrix, that is, $A=(a(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M},B=(b(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M}$. The eigenvalue problem (1) can be formulated as the following algebraic eigenvalue problem:
$A\bm{w}=\lambda B\bm{w},$ (3)
where $\bm{w}=(w^{(1)},\cdots,w^{(M)})^{T}\in\mathbb{R}^{M}.$
By solving the algebraic eigenvalue problem (3), an approximate solution to the eigenvalue problem (1) is then obtained.
@@ 2.2 Subspace method based on neural networks
In [57], the following general source problem is considered:
$\begin{cases}\mathscr{A}u=f&\text{in }\rm\Omega,\\ \mathscr{B}u=g&\text{on }\partial\rm\Omega,\end{cases}$
where $\mathscr{A}$ and $\mathscr{B}$ are two differential operators, $f$ and $g$ are given functions.
The neural networks architecture used in [57] is comprised of four parts: an input layer, several hidden layers, a subspace layer and an output layer. For simplicity, we take the case of single output as an example to show the mathematical expression of the neural networks. The expression of multiple outputs can be obtained similarly.
Denote $L$ the number of hidden layers, $n_{0}=d$ the number of neurons in the input layer, $n_{1},\cdots,n_{L}$ the number of neurons in each hidden layer, $n_{L+1}=M$ the number of neurons in the subspace layer. The neural networks used in [57] can be expressed as follows:
@@
Here, $\mathcal{T}_{l}(\bm{y})=W_{l}\cdot\bm{y}+\bm{b}_{l}$, $W_{l}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $\bm{b}_{l}\in\mathbb{R}^{n_{l}}$ are the weights and biases of the corresponding layer, respectively, $\bm{x}\in\mathbb{R}^{d}$ is the input, $\bm{\varphi}=(\varphi_{1},\cdots,\varphi_{M})^{T}$ and $\bm{w}=(w^{(1)},\cdots,w^{(M)})^{T}$, $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is an activation function which can be extended to $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ as follows: $\sigma(\bm{z})=[\sigma(z_{1}),\cdots,\sigma(z_{n})]^{T}\text{for any}\ \bm{z}=(z_{1},\cdots,z_{n})^{T}\ \in\mathbb{R}^{n}.$ Denote $\theta=\{W_{1},\cdots,W_{L+1},\bm{b}_{1},\cdots,\bm{b}_{L+1}\}\in{\rm\Theta}:=\mathbb{R}^{n_{1}\times n_{0}}\times\cdots\times\mathbb{R}^{n_{L+1}\times n_{L}}\times\mathbb{R}^{n_{1}}\times\cdots\times\mathbb{R}^{n_{L+1}}$ the parameters of the neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$, and $u(\bm{x};\theta,\bm{w})=\sum\limits_{i=1}^{M}w^{(i)}\varphi_{i}(\bm{x};\theta)$ are the output of neural networks with respect to the input $\bm{x}$ and parameters $\theta$ and $\bm{w}$.
Next, we will review the general framework of the subspace method based on neural networks for the partial differential equations proposed in [57]. Firstly, the neural networks architecture introduced above is adopted, and parameters $\theta$ and $\bm{w}$ are initialized. Next, train the neural network-based functions, i.e., update parameters $\theta$ by minimizing some loss function, aiming to obtain functions that are better than the initial ones through training. After training, find an approximate solution in the subspace spanned by the trained neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$, i.e., update parameters $\bm{w}$ by solving some linear system, aiming to obtain an approximate solution of the partial differential equation. The above process is summarized as a general framework for solving the partial differential equations in [57]. We also list it here for the completeness, see Algorithm 1 for the details.
Algorithm 1 General framework of the subspace method based on neural networks for partial differential equations
 1. Initialize the neural network architecture, give and then fix $\bm{w}_{0}$.
 2. Define some loss function $\mathcal{L}(\theta,\bm{w}_{0})$ using information contained in the PDEs, then obtain parameters $\theta$ by minimizing the loss function using some optimization method
$\theta^{\ast}=\mathop{\arg\min}\limits_{\theta\in\rm\Theta}\mathcal{L}(\theta,\bm{w}_{0}),$
and obtain the updated neural network-based functions $\{\varphi_{i}(\bm{x};\theta^{\ast})\}_{i=1}^{M}$
 3. Using some classical method to find the approximation for the solution of the original problems in the trained subspace spanned by $\{\varphi_{i}(\bm{x};\theta^{\ast})\}_{i=1}^{M}$, i.e., obtain optimal $\bm{w}^{\ast}$, and then obtain the approximate solution $u(\bm{x};\theta^{\ast},\bm{w}^{\ast})$.